OpenStudy (ashleynguyenx3):

Probability: How many different permutations are there if you used 4 different digits? Repeat a digit?

OpenStudy (unknownunknown):

4, 4^2, 4^3, 4^n.... n=repeats

OpenStudy (ashleynguyenx3):

Wouldn't it be 4 factorial?

OpenStudy (ashleynguyenx3):

Or is it different?

OpenStudy (unknownunknown):

That's if you want unique digits, then yes.

OpenStudy (unknownunknown):

Ahh I see what you mean. You take one out each time? So 4.. 3.. 2.. 1.. yes that's 4!.

OpenStudy (ashleynguyenx3):

Would it still be that if order matters?

OpenStudy (unknownunknown):

If you want certain digits in a specific order, then that would change it.

OpenStudy (unknownunknown):

Probability of selecting one digit from 4 is 0.25. Then 1 digit from 3 is 0.33, 1 digit from 2 is 0.5, 1 digit from 1 is 1. So the probability of doing all of that correctly is 0.25*0.33*0.5

OpenStudy (ashleynguyenx3):

I think I'm supposed to find the number or permutations not the probability, so I'm not completely sure.

OpenStudy (unknownunknown):

Then I'd go with 4!. Order wouldn't matter there.

OpenStudy (ashleynguyenx3):

Okay, so what about when we repeat a number?

OpenStudy (unknownunknown):

If we repeat, then instead of doing 4 * 3 * 2 *1, we have four digits still in each case. So we do 4 * 4 * 4 * 4 = 4^4 permutation

OpenStudy (unknownunknown):

By repeat, you mean we don't exclude the number we pick, right?

OpenStudy (ashleynguyenx3):

No this question is more for like iphone security codes. If that makes sense.

OpenStudy (ashleynguyenx3):

So that's why order matters.

OpenStudy (unknownunknown):

Okay, can you give me an example of two combinations then, that follow what you require.

OpenStudy (ashleynguyenx3):

So for 3 digits it would be like 123, 132, 213, 231, 321, 312 if there is no repeating

OpenStudy (ashleynguyenx3):

With a repeating it would be 112, 121, 211

OpenStudy (ashleynguyenx3):

I know the one without repeating is 4!

OpenStudy (unknownunknown):

So for any two numbers a and b, a appears twice, and b once, in all three combinations. And do that for all numbers from 1 to 4?

OpenStudy (ashleynguyenx3):

Yeah I think I found a site that might help

OpenStudy (unknownunknown):

Alright, good luck.